Determine the [H+] , pH, and pOH of a solution with an [OH−] of 7.2×10−8 M at 25 °C.

[H+]= 

pH=

pOH=



1
Expert's answer
2022-04-24T18:58:05-0400

pOH = -lg[OH−] = -lg(7.2×10-8) = 7.14

pH = 14 - pOH = 14 - 7.14 = 6.86

[H+] = 10-pH = 10-6.86 = 1.38×10-7 M